Plz explain why. Thank you
6. E Go A plate carries a charge of -3.0 HC, while a rod carries a charge of +2.0 uC. How many electrons must be transferred from the plate to the rod, so that both objects have the same charge? 7. -HINT Water has a mass per mole of 18.0 g/mol, and each water mo- lecule (H,0) has 10 electrons. (a) How many electrons are there in one liter (1.00x 10-m) of water? (b) What is the net charge of all these electrons?
Answer #1

Total Charge = (-3 + 2) = -1 mu c

both object have a charge of -0.50mucmu c so   -2.50mucmu c charge will transfer from -3.0 mu c .

Number of electron transfer :-

q = ne

n= number of electron

e= electronic charge

  2.50 10-6 1.6 10-191.6 10-13 rl

Part-b

Avogadro's number is the quantity of representative particles in a mole of a substance. This value is 6.02 x 1023 representative particles per mole. A representative particle is an atom, molecule or electron, to name a few.

(a) 1000q 1mol 6.02310molecules 10electron 23 IL * L 18g lmol = 3.34*1026electron lmolecule

Total Char ge 3.34 1025 electrons * = 5.344 * 108columb 1.6 101 Columl electron